OpenStudy (anonymous):

If sin x = .42, what is cos x?

OpenStudy (anonymous):

??

OpenStudy (campbell_st):

well draw a triangle 0.42 = 42/100  find b using pythagoras then frind the cos(x)

OpenStudy (anonymous):

or compute \[\large \pm\sqrt{1-.42^2}\]

OpenStudy (anonymous):

I dont understand @campbell_st

OpenStudy (campbell_st):

well do you know pythagoras' theorem... c^2 = a^2 + b^2 it can be used the find the measure of b this is because you are given the value of sin(x)... which is opposite/hypotenuse so you have the measures for the hypotenuse and opposite side. You just need to find the adjacent side before find cos(x) b^2 = 100^2 - 42^2 so find b then cos(x) = b/100

OpenStudy (anonymous):

If tan x = .76 and cos x = .22, what is sin x? @campbell_st

OpenStudy (campbell_st):

well you know that  hope it helps

OpenStudy (anonymous):

Thanks, but how would you do this one If cos x = .75, what is sin x? @campbell_st